OpenStudy (babtaooche):

Help! I'll fan and medal Find the sum of the summation of 3 i plus2, from i equals 5 to 13

OpenStudy (plainntall):

You mean \[\sum_{5}^{13}3i+2\]

OpenStudy (babtaooche):

yes except where the 5 is, it says i=5

OpenStudy (zzr0ck3r):

Hint: \(\sum_{i=5}^{13}3i+2=3\sum_{i=5}^{13}i+\sum_{i=5}^{13}2=3\sum_{i=5}^{13}i+2\sum_{i=5}^{13}1\)

OpenStudy (babtaooche):

These are really confusing me for and I'm not sure how to do them

OpenStudy (threshsupport):

mods? i need help it won't let me post a question without paying

OpenStudy (zzr0ck3r):

Do you know how to find \(\sum_{i=5}^{13}i\) ?

OpenStudy (babtaooche):

No

OpenStudy (babtaooche):

Yes, It's 9i

OpenStudy (zzr0ck3r):

\(\sum_{i=1}^{n}i=\dfrac{n(n+1)}{2}\) So we have \(\sum_{i=5}^{13}i=\sum_{i=1}^{13}i-\sum_{i=5}^{5}=\dfrac{13*14}{2}-\dfrac{5*6}{2}\)

OpenStudy (zzr0ck3r):

that should say \(\sum_{i=5}^{13}i=\sum_{i=1}^{13}i-\sum_{i=4}^{5}=\dfrac{13*14}{2}-\dfrac{4*5}{2}\)

OpenStudy (babtaooche):

That makes more sense

OpenStudy (zzr0ck3r):

so \(\sum_{i=5}^{13}3i+2=3\sum_{i=5}^{13}i+\sum_{i=5}^{13}2=3\sum_{i=5}^{13}i+2\sum_{i=5}^{13}1\\=3(\dfrac{13*14}{2}-\dfrac{4*5}{2})-2(9)\)

OpenStudy (babtaooche):

Alright, I'm gonna see if I can solve it

OpenStudy (zzr0ck3r):

got to go, good luck. Ill check back later.